# Quantum Key Leasing for PKE and FHE with a Classical Lessor

### 03 Jan 2024 v3

@@ 2 Technical Overview
While the scheme above gives rise to a PKE scheme satisfying a weak but a non-trivial security guarantee, such a definition is not really enough. It might be possible that an adversary simply guesses the correct certificate $(\alpha,\mathbf{d})$ by picking $\mathbf{d}$ randomly and choosing $\alpha$ at random as well. With $\frac{1}{2}$ probability, this adversary might be right in producing a valid certificate. Since its state is still preserved, it could continue to successfully decrypt all ciphertexts (moreover, it could simply measure the state in the standard basis and keep a classical key). We’d like to achieve a stronger definition where even if the certificate passes with a noticable probability, upon successful passing, the advantage should not be non-negligible.
Our main idea there is that one could create $\lambda$ independent instances of the scheme where one encrypts secret sharing of the bit $\mu$ where $\lambda$ is the security parameter. The idea is that now the public key would consist of $\lambda$ indepdent matrices $\mathbf{A}_{i}$, $\lambda$ indepdent binary secret LWE samples $\mathbf{k}_{i}=\mathbf{s}_{i}\mathbf{A}_{i}+\mathbf{e}_{i}$ along with $\mathbf{y}_{i}=\mathbf{x}_{i,0}\mathbf{A}_{i}+\mathbf{e}^{\prime}_{i}$ for $i\in[\lambda]$. The deletion certificate would consist of $\lambda$ vectors $(\alpha_{i},\mathbf{d}_{i})$ and each of them are indepdently verified. The hope again is that if one is able to distinguish encryptions of 00 from encryptions of $1$, then such an adversary should essentially be able to distinguish for each component $i\in[\lambda]$. We could use this to exctract all $\{\mathbf{x}_{i,b_{i}}\}_{i\in[\lambda]}$($b_{i}=0$ or 1). Then, we can appeal to the results from the parallel repetition variant of the game used for proofs of quantumness. This variant has been studied in [RS19]. The soundness property ensures that it is computationally hard to come up with noticable probability valid responses $\{(\alpha_{i},\mathbf{d}_{i})\}_{i\in[\lambda]}$ and $\{\mathbf{x}_{i,b_{i}}\}_{i\in[\lambda]}$($b_{i}=0$ or 1) for all indices.
@@
After we modify the encryption algorithm to compute $\mathbf{B}\mathbf{R}+\mathbf{E}+\mu\mathbf{G}$ where $\mathbf{E}$ is zero everywhere else except the last row containing discrete Gaussian with parameter $\sigma^{\prime\prime}$ superpolnomially more than the norm of $\mathbf{e}^{\prime}_{i}\mathbf{R}$, we make a nice observation: the last row of $\mathbf{B}\mathbf{R}+\mathbf{E}$ would now be distributed as an LWE sample: in terms of the secret $(\mathbf{x}_{1,b_{1}},\mathbf{x}_{2,b_{2}},\ldots,\mathbf{x}_{\lambda,b_{\lambda}})$($b_{i}=0\text{ or }1$) with the coefficient vector $[\mathbf{V}^{\top}_{1},\ldots,\mathbf{V}^{\top}_{\lambda}]^{\top}$. Now we could appeal the LHL to replace $\{(\mathbf{A}_{i}\mathbf{R},\mathbf{V}_{i}\mathbf{R},\mathbf{k}_{i}\mathbf{R})\}_{i\in[\lambda]}$ to completely random and the last row by a fresh sample with error according to discrete Gaussian in parameter $\sigma^{\prime\prime}$, with the long secret $((\mathbf{x}_{1,b_{1}},\mathbf{x}_{2,b_{2}},\ldots,\mathbf{x}_{\lambda,b_{\lambda}})$ and random coefficient vector $[\mathbf{V}^{\top}_{1},\ldots,\mathbf{V}^{\top}_{\lambda}]^{\top}$. If an adversary now distinguish such ciphertexts from random, we should be able to extract the entire long secret vector in one shot $(\mathbf{x}_{1,b_{1}},\mathbf{x}_{2,b_{2}},\ldots,\mathbf{x}_{\lambda,b_{\lambda}})$ using our proposed quantum search-to-decision reduction.
@@
Proof Outline
We now go into slight more details about our high level proof strategy. Recall that in the security game, a successful attacker would first need to output a valid deletion certifiacte (which we denote by the event $\mathsf{CertPass}$) along with that it must produce a state $\rho_{\mathsf{delete}}$ for which our test of good decryptor $\mathsf{TI}_{\frac{1}{2}+\gamma}$ for some noticable $\gamma$ passes with invere-polynomial probability. Namely $\mathsf{TI}_{\frac{1}{2}+\gamma}(\rho_{\mathsf{delete}})=1$ with inverse polynomial probability. We call this event $\mathsf{GoodDecryptor}$. It must hold that $\Pr[\mathsf{GoodDecryptor}\wedge\mathsf{CertPass}]$ is noticable for such successful attacker. Since we are guaranteed that there is a noticable chance of overlap of the two events, our hope is that our extraction procedure would extract $\{\mathbf{x}_{0,i}\}$ just with enough probability to induce a noticable probability overlap between $\mathsf{CertPass}$ and successful extraction causing our reduction to win the parallel repetion game of Radian et. al. We call the event when the extraction holds as $\mathsf{ExtractionOccurs}$.
At this point it is tempting to think of a search-to-decision classical reduction for LWE and compile it to a quantum-reduction via known menthods such as the one by Bitansky et. al. [BBK22]. Unfortunately, we don’t have a citable theorem from this work to use directly [^4] As discussed in Section 1.2, in fact all existing works on post-quantum LWE search-to-decision reductions are not directly applicable to our setting.. We need precise bounds on the probability of success, and the running time for the extraction.
Let us briefly explain the story: for the extraction to occur we would need to move to a world where $\mathbf{k}_{i}$’s are switched with random. While this won’t change the probability of $\mathsf{GoodDecryptor}$ by a non-negligible amount due to LWE security, in this hybrid deletion certificates will no longer exist. We could infer from there using reductions compiled using [BBK22]) that $\mathsf{ExtractionOccurs}$ with probability $\frac{1}{{\sf poly}(\lambda)}$ for some polynomial ${\sf poly}$. This probability of extraction will still be the same (up to a negligible loss) if we went back to the world where $\{\mathbf{k}_{i}\}$ are again from the LWE distribution. However, we can’t infer from this that the event of $\mathsf{ExtractionOccurs}$ overlaps with $\mathsf{CertPass}$. A similar issue was encountered by Ananth et. al. [APV23]. To address this issue the [APV23] had to rely on new conjectures.
To address this issue, we devise a high probability search-to-decision reduction that extracts (in the world where $\mathbf{k}_{i}$’s are random) with probability $1-\mathsf{negl}(\lambda)$ whenever $\mathsf{GoodDecryptor}$ holds. Thus when we switch $\mathbf{k}_{i}$’s with LWE, due to LWE security, $\mathsf{ExtractionOccurs}$ also succeeds with all but negligible security whenever $\mathsf{GoodDecryptor}$ holds. Since we are guaranteed that there is a noticeable overlap between $\mathsf{GoodDecryptor}$ and $\mathsf{CertPass}$, this implies a noticeable overlap between $\mathsf{CertPass}$ and $\mathsf{ExtractionOccurs}$.
@@
    For a successful adversary in our game, it must produce a decryptor $\rho$ that is good with probability $p$ for some noticeable $p$ (apart from giving a valid certificate with oticeable probability).
@@
Our quantum search-to-decision reduction the quanis inspired by Regev’s search to decision reduction [Reg05b]. We now describe the setup of our reduction. We consider a simpler setup than in our technical section (which is a bit specific to our setting).
@@
We first describe the classical intuition and then describe how to lift that intution to the quantum setting. Classically, we can consider recovering $\mathsf{x}$ coordinate by coordinate. Say the first coordinate $x_{1}\in[-B,B]$, we could choose a total of $2B$ guesses. For each guess $g$, we could consider the process of generating samples as:
@@
If our guess was correct, the sample that we end up producing $(\mathbf{A}^{\prime},\mathbf{b}^{\prime})$ is distributed statistically close to the distribution for the distinguishing problem (LWE with secret $\mathbf{x}$). This is because $\mathbf{e}^{\prime}+\mathbf{e}\mathbf{R}$ due to noise flodding is within ${\sf poly}(m)\cdot\frac{\sigma}{\sigma^{\prime}}$ statistical distance. Similarly, $\mathbf{A}^{\prime}$ due to LHL is within $q^{-n}$ statistical distance from uniform if $m$ is sufficiently large. This means that the distribution is within $\eta={\sf poly}(m)\frac{\sigma}{\sigma^{\prime}}$ statistical distance from the relevant distinguishing problem.
On the other hand, if the guess is incorrect, then one can observe that the distribution produces samples that are within similar statistical distance $\eta={\sf poly}(m)\cdot\frac{\sigma}{\sigma^{\prime}}$ from truly random distribution.
Thus, if the guess is correct, a classical adversary can distinguish the above distribution from random with probability at least $\frac{1}{2}+\gamma-O(\eta)$, and likewise if the guess is incorrect the maximum distinguishing probability is $\frac{1}{2}+O(\eta)$. We could therefore test the adversary by making roughly ${\sf poly}(\frac{1}{\gamma},\log\delta)$ calls to the distinguisher to guarantee the guess is correct with probability $1-\delta$. Setting $\delta=p^{-n}$, the reduction will extract $\mathbf{x}$ in time that’s polynomial in $B,n,m,\log p,\frac{1}{\gamma}$ (bound on $\mathbf{x}$’s norm).
@@
Moving to the quantum setting, we have to address a number of challenges. If we use our state to distinguish and LWE sample from random in the way above, the state could get destructed preventing us from doing repetitions. In particular, the classical reduction needs to run the distinguisher many times over randomized inputs and "measure" its outcome to obtain useful information, which seems implausible when using a quantum distinguisher.
To overcome these issue, we will leverage the properties of Threshold Implementation. We make a key observation that the procedure where we repeatedly create samples according to our guess $g$ and check on the distinguisher’s output distribution to get an estimate on whether $g$ is correct, can happen "inside" the measurement procedure ${\sf TI}$.
@@
  • ${\sf TI}_{\sf unif}$ projects onto distinguishers good at distinguishing uniform random samples from uniform random samples with probability at least $1/2+\gamma$.
@@
  • As we have discussed in the classical setting, when the guess $g$ is correct, we get to create samples statistically close to LWE samples; when the guess $g$ is incorrect, we get to create samples close to uniform random. Combining these with properties of ${\sf TI}_{g}$ for distributionally close measurements, we can show that:
@@
    • When the guess is incorrect: If we apply projection ${\sf TI}_{g}$ on it for an incorrect guess $g$, we will have $\Pr[(\mathbf{I}-{\sf TI}_{g})(\rho^{\prime})=1]=\Pr[{\sf TI}_{g}(\rho^{\prime})=0]$ equal to $1$. In this case, we are asking the distinguisher to distinguish between uniform vs. negligibly-close-uniform. Clearly, no distinguisher can distinguish with noticeable advantage $\gamma$ if the statistical distance is negligible. Therefore ${\sf TI}_{\sf unif}$ will output 0 for all possible states.
@@
A key observation is that, in both cases above, because we obtain a certain measurement outcome with probability 1, the leftover state $\rho^{\prime}$ will remain unchanged in terms of its usability: in the case where ${\sf TI}_{g}$ outputs 1, we know that if we apply ${\sf TI}_{g}$ for a correct $g$ again, we get the same outcome with probability $1$ due to the projective property; in the case where ${\sf TI}_{g}$ outputs 0, since ${\sf TI}_{g}$ will always output 0 on any $\rho$ with probability 1, the state will remain undisturbed. Thus, we could find out $\mathbf{x}$ entry by entry while keeping our quantum state’s performance almost unchanged.
Actual Algorithm: Approximate Threshold Implementation
@@
However, ${\sf TI}$ is not an efficient measurement and we need to make the above procedure efficient. Fortunately, the work of Zhandry [Zha20], showed how to replace ${\sf TI}$ by an efficient, approximate projection, namely approximate threshold implementation(${\sf ATI}$). These approximate projections induce an error upon each measurement, but these errors are easy to account for by the properties of these approximate projections and our choice of parameters.
@@
In the case when the approximate threshold implementation ${\sf ATI}_{g}$ corresponds to a correct guess $g$, we use properties similar to those shown in [Zha20] to demonstrate that the next ${\sf ATI}_{g}$ with a correct guess will give the same outcome with probability close to 1.
@@
The remaining question is: what happens when we apply ${\sf ATI}_{g}$ with an incorrect $g$? Intriguingly, we make the following new observation (see Section 5.2.1): the operation ${\sf ATI}_{g}$ with an incorrect $g$ acts almost like an identity operator on any state and hardly changes the state itself. Therefore, such measurements incur only negligible deviation additively on the measurement outcomes when it comes to the next ${\sf ATI}_{g}$ with correct $g$’s.
In the end, we further materialize our approach with some proper choice of parameters so that the errors will not accumulate on distinguisher’s state to a degree that affects the algorithm’s performance. Conclusively, the exact reduction in our scheme works if the number of guesses are superpolynomially lesser than inverse the statistical distance. See details in Section 11. On the other hand, if we consider a clean setting where our input distributions are real LWE versus truly uniform, our reduction works for any subexponential number of guesses (Appendix D), if working with subexponential-hardness of LWE.

### 06 Mar 2024 v4

@@ 2 Technical Overview
While the scheme above gives rise to a PKE scheme satisfying a weak but a non-trivial security guarantee, such a definition is not really enough. It might be possible that an adversary simply guesses the correct certificate $(\alpha,\mathbf{d})$ by picking $\mathbf{d}$ randomly and choosing $\alpha$ at random as well. With $\frac{1}{2}$ probability, this adversary might be right in producing a valid certificate. Since its state is still preserved, it could continue to successfully decrypt all ciphertexts (moreover, it could simply measure the state in the standard basis and keep a classical key). We’d like to achieve a stronger definition where even if the certificate passes with a noticeable probability, upon successful passing, the advantage should not be non-negligible.
Our main idea there is that one could create $\lambda$ independent instances of the scheme where one encrypts secret sharing of the bit $\mu$ where $\lambda$ is the security parameter. The idea is that now the public key would consist of $\lambda$ independent matrices $\mathbf{A}_{i}$, $\lambda$ independent binary secret LWE samples $\mathbf{k}_{i}=\mathbf{s}_{i}\mathbf{A}_{i}+\mathbf{e}_{i}$ along with $\mathbf{y}_{i}=\mathbf{x}_{i,0}\mathbf{A}_{i}+\mathbf{e}^{\prime}_{i}$ for $i\in[\lambda]$. The deletion certificate would consist of $\lambda$ vectors $(\alpha_{i},\mathbf{d}_{i})$ and each of them are independently verified. The hope again is that if one is able to distinguish encryptions of 00 from encryptions of $1$, then such an adversary should essentially be able to distinguish for each component $i\in[\lambda]$. We could use this to extract all $\{\mathbf{x}_{i,b_{i}}\}_{i\in[\lambda]}$($b_{i}=0$ or 1). Then, we can appeal to the results from the parallel repetition variant of the game used for proofs of quantumness. This variant has been studied in [RS19]. The soundness property ensures that it is computationally hard to come up with noticable probability valid responses $\{(\alpha_{i},\mathbf{d}_{i})\}_{i\in[\lambda]}$ and $\{\mathbf{x}_{i,b_{i}}\}_{i\in[\lambda]}$($b_{i}=0$ or 1) for all indices.
@@
After we modify the encryption algorithm to compute $\mathbf{B}\mathbf{R}+\mathbf{E}+\mu\mathbf{G}$ where $\mathbf{E}$ is zero everywhere else except the last row containing discrete Gaussian with parameter $\sigma^{\prime\prime}$ superpolynomially more than the norm of $\mathbf{e}^{\prime}_{i}\mathbf{R}$, we make a nice observation: the last row of $\mathbf{B}\mathbf{R}+\mathbf{E}$ would now be distributed as an LWE sample: in terms of the secret $(\mathbf{x}_{1,b_{1}},\mathbf{x}_{2,b_{2}},\ldots,\mathbf{x}_{\lambda,b_{\lambda}})$($b_{i}=0\text{ or }1$) with the coefficient vector $[\mathbf{V}^{\top}_{1},\ldots,\mathbf{V}^{\top}_{\lambda}]^{\top}$. Now we could appeal the LHL to replace $\{(\mathbf{A}_{i}\mathbf{R},\mathbf{V}_{i}\mathbf{R},\mathbf{k}_{i}\mathbf{R})\}_{i\in[\lambda]}$ to completely random and the last row by a fresh sample with error according to discrete Gaussian in parameter $\sigma^{\prime\prime}$, with the long secret $(\mathbf{x}_{1,b_{1}},\mathbf{x}_{2,b_{2}},\ldots,\mathbf{x}_{\lambda,b_{\lambda}})$ and random coefficient vector $[\mathbf{V}^{\top}_{1},\ldots,\mathbf{V}^{\top}_{\lambda}]^{\top}$. If an adversary now distinguish such ciphertexts from random, we should be able to extract the entire long secret vector in one shot $(\mathbf{x}_{1,b_{1}},\mathbf{x}_{2,b_{2}},\ldots,\mathbf{x}_{\lambda,b_{\lambda}})$ using our proposed quantum search-to-decision reduction.
@@
    For a successful adversary in our game, it must produce a decryptor $\rho$ that is good with probability $p$ for some noticeable $p$ (apart from giving a valid certificate with noticeable probability).
@@
Proof Outline
We now go into more details about our high level proof strategy. Recall that in the security game, a successful attacker would first need to output a valid deletion certificate (which we denote by the event $\mathsf{CertPass}$) along with that it must produce a state $\rho_{\mathsf{delete}}$ for which our test of good decryptor $\mathsf{TI}_{\frac{1}{2}+\gamma}$ for some noticeable $\gamma$ passes with inverse-polynomial probability. Namely, $\mathsf{TI}_{\frac{1}{2}+\gamma}(\rho_{\mathsf{delete}})=1$ with inverse polynomial probability. We call this event $\mathsf{GoodDecryptor}$. It must hold that $\Pr[\mathsf{GoodDecryptor}\wedge\mathsf{CertPass}]$ is noticeable for such a successful attacker. Since we are guaranteed that there is a noticeable chance of overlap of the two events, our hope is that our extraction procedure would extract $\{\mathbf{x}_{0,i}\}$ just with enough probability to induce a noticeable probability overlap between $\mathsf{CertPass}$ and a successful extraction causing our reduction to win the parallel repetition game of Radian et al. We call the event when the extraction holds as $\mathsf{ExtractionOccurs}$.
At this point it is tempting to think of a search-to-decision classical reduction for LWE and compile it to a quantum-reduction via known methods such as the one by Bitansky et al. [BBK22]. Unfortunately, we don’t have a citable theorem from this work to use directly [^5] As discussed in Section 1.2, in fact all existing works on post-quantum LWE search-to-decision reductions are not directly applicable to our setting.. We need precise bounds on the probability of success, and the running time for the extraction.
Let us briefly explain the story: for the extraction to occur, we would need to move to a world where $\mathbf{k}_{i}$’s are switched with random. While this won’t change the probability of $\mathsf{GoodDecryptor}$ by a non-negligible amount due to LWE security, in this hybrid deletion certificates will no longer exist. We could infer from there using reductions compiled using [BBK22]) that $\mathsf{ExtractionOccurs}$ with probability $\frac{1}{{\sf poly}(\lambda)}$ for some polynomial ${\sf poly}$. This probability of extraction will still be the same (up to a negligible loss) if we went back to the world where $\{\mathbf{k}_{i}\}$ are again from the LWE distribution. However, we can’t infer from this that the event of $\mathsf{ExtractionOccurs}$ overlaps with $\mathsf{CertPass}$. A similar issue was encountered by Ananth et. al. [APV23]. To address this issue the [APV23] had to rely on new conjectures.
To address this issue, we devise a high probability search-to-decision reduction that extracts (in the world where $\mathbf{k}_{i}$’s are random) with probability $1-\mathsf{negl}(\lambda)$ whenever $\mathsf{GoodDecryptor}$ holds. Thus, when we switch $\mathbf{k}_{i}$’s with LWE, due to LWE security, $\mathsf{ExtractionOccurs}$ also succeeds with all but negligible security whenever $\mathsf{GoodDecryptor}$ holds. Since we are guaranteed that there is a noticeable overlap between $\mathsf{GoodDecryptor}$ and $\mathsf{CertPass}$, this implies a noticeable overlap between $\mathsf{CertPass}$ and $\mathsf{ExtractionOccurs}$.
@@
Our quantum search-to-decision reduction is inspired by Regev’s search to decision reduction [Reg05b]. We now describe the setup of our reduction. We consider a simpler setup than in our technical section (which is a bit specific to our setting).
@@
We first describe the classical intuition and then describe how to lift that intuition to the quantum setting. Classically, we can consider recovering $\mathbf{x}$ coordinate by coordinate. Say the first coordinate $x_{1}\in[-B,B]$, we could choose a total of $2B$ guesses. For each guess $g$, we could consider the process of generating samples as:
@@
If our guess was correct, the sample that we end up producing $(\mathbf{A}^{\prime},\mathbf{b}^{\prime})$ is distributed statistically close to the distribution for the distinguishing problem (LWE with secret $\mathbf{x}$). This is because $\mathbf{e}^{\prime}+\mathbf{e}\mathbf{R}$ due to noise flooding is within ${\sf poly}(m)\cdot\frac{\sigma}{\sigma^{\prime}}$ statistical distance. Similarly, $\mathbf{A}^{\prime}$ due to LHL is within $p^{-n}$ statistical distance from uniform if $m$ is sufficiently large. This means that the distribution is within $\eta={\sf poly}(m)\frac{\sigma}{\sigma^{\prime}}$ statistical distance from the relevant distinguishing problem for a corect guess, i.e. the "LWE" case.
On the other hand, if the guess is incorrect, then one can observe that the distribution produces samples that are within even smaller statistical distance $\eta^{\prime}={\sf poly}(m)\cdot\frac{1}{p^{n}}$ from truly random distribution (no noise flooding required).
Thus, if the guess is correct, a classical adversary can distinguish the above distribution from random with probability at least $\frac{1}{2}+\gamma-O(\eta)$, and likewise if the guess is incorrect the maximum distinguishing probability is $\frac{1}{2}+O(\eta^{\prime})$. We could therefore test the adversary by making roughly ${\sf poly}(\frac{1}{\gamma},\log\delta)$ calls to the distinguisher to guarantee the guess is correct with probability $1-\delta$. Setting $\delta=p^{-n}$, the reduction will extract $\mathbf{x}$ in time that’s polynomial in $B,n,m,\log p,\frac{1}{\gamma}$ (bound on $\mathbf{x}$’s norm).
@@
Moving to the quantum setting, we have to address a number of challenges. If we use our state to distinguish and LWE sample from random in the way above, the state could get destructed, preventing us from doing repetitions. In particular, the classical reduction needs to run the distinguisher many times over randomized inputs and "measure" its outcome to obtain useful information, which seems implausible when using a quantum distinguisher.
To overcome these issue, we will leverage the properties of Threshold Implementation. We make a key observation that the procedure, where we repeatedly create samples according to our guess $g$ and check on the distinguisher’s output distribution to get an estimate on whether $g$ is correct, can happen "inside" the measurement procedure ${\sf TI}$.
@@
  • ${\sf TI}_{\sf unif}$ projects onto distinguishers good at distinguishing close-to-uniform random samples from uniform random samples with probability at least $1/2+\gamma$.
@@
  • As we have discussed in the classical setting, when the guess $g$ is correct, we get to create samples statistically close to LWE samples; when the guess $g$ is incorrect, we get to create samples close to uniform random. Combining these with properties of ${\sf TI}_{g}$ for statistically close measurements, we can show that:
@@
    • When the guess is incorrect: If we apply projection ${\sf TI}_{g}$ on it for an incorrect guess $g$, we will have $\Pr[(\mathbf{I}-{\sf TI}_{g})(\rho^{\prime})=1]=\Pr[{\sf TI}_{g}(\rho^{\prime})=0]$ equal to $1-\mathsf{negl}(\lambda)$. In this case, we are asking the distinguisher to distinguish between uniform vs. negligibly-close-uniform. Clearly, no distinguisher can distinguish with noticeable advantage $\gamma$ if the statistical distance is negligible. Therefore, ${\sf TI}_{\sf unif}$ will output 0 for all possible states.
@@
A key observation is that, in both cases above, because of the measurement’s projective property , the leftover state $\rho^{\prime}$ will remain unchanged in terms of its usability: in the case where ${\sf TI}_{g}$ outputs 1, we know that if we apply ${\sf TI}_{g}$ for a correct $g$ again, we get the same outcome with probability $1$ due to the projective property; in the case of incorrect $g$’s, similarly, after we have the first ${\sf TI}_{g}$ that outputs 0, the remaining ${\sf TI}_{g}$ will always output 0 on any $\rho$ with probability 1. Therefore, the state will remain undisturbed. Thus, we could find out $\mathbf{x}$ entry by entry while keeping our quantum state’s performance almost unchanged.
Actual Algorithm: New Observations on the Approximate Threshold Implementation
@@
However, the ${\sf TI}$ used above is not an efficient measurement and we need to make the above procedure efficient. Fortunately, the work of Zhandry [Zha20], showed how to replace ${\sf TI}$ by an efficient, approximate projection, namely approximate threshold implementation(${\sf ATI}$). These approximate projections induce an error upon each measurement, and since they are not "exact projections", we cannot make the same argument as in the above ${\sf TI}$ analysis. We account for these issues by using the properties of these approximate projections.
@@
In the case when the approximate threshold implementation ${\sf ATI}_{g}$ corresponds to a correct guess $g$, we use properties similar to those shown in [Zha20] to demonstrate that the next ${\sf ATI}_{g}$ with a correct guess will give the same outcome with probability close to 1, similar to the clean ${\sf TI}$ setting.
@@
The remaining question is: what happens when we apply ${\sf ATI}_{g}$ with an incorrect $g$? Intriguingly, we make the following new observation (see details in Section 5.2.1): the operation ${\sf ATI}_{g}$ with an incorrect $g$ acts almost like an identity operator on any state and hardly changes the state itself, in terms of trace distance. When we use the [Zha20, MW05] ${\sf ATI}$ measurement to test if a quantum adversary can distinguish between two identical distributions, naturally the measurement will output 00 with probability close to 1 (because no adversary can possibly distinguish). More importantly, the measurement acts exactly like an identity operator. When we test a quantum adversary on distinguishing two close statistically-close distributions, the ${\sf ATI}$ measurement acts close to an identity operator. In a nutshell, each such ${\sf ATI}_{g}$ with incorrect $g$ incurs only exponentially small deviation additively on the trace distance of the state. Therefore, after we have applied many ${\sf ATI}_{g}$ with incorrect $g$’s, our next ${\sf ATI}_{g}$ with correct $g$ will yield outcome 1 with overwhelming probability as the previous ${\sf ATI}_{g}$ with correct $g$ does because the state is almost unchanged since that time.
In the end, we make sure that the measurement errors will not accumulate on distinguisher’s state to a degree that affects the algorithm’s performance. Conclusively, the dominating error is incurred by every measurement with ${\sf ATI}_{g}$ with a correct $g$, which is some inverse superpolynomial. But this error will only accumulate additively for polynomially many times (the number of measurements with correct $g$ is the same as the dimension of the LWE secret). On the other hand, the errors incurred by ${\sf ATI}_{g}$ with an incorrect $g$ are exponentially small. Therefore, the overall loss is negligible. The exact reduction used in our security proof works for any subexponential number of guesses, when working with subexponential-hardness of LWE. See details in Section 11.
More generically, if we consider a clean setting where our input distributions are real LWE versus truly uniform samples, we even have less to worry about because all losses in the above measurements will be exponentially small. Therefore, our search-to-decision reduction works for any LWE parameters (Appendix D) and runs in time roughly ${\sf poly}(\lambda,n,B)$ where $n$ is the dimension of the LWE secret and $B$ is the domain size of each entry of the secret.